OpenStudy (anonymous):

The weight of a liquid varies directly as its volume V. If the weight of the liquid in a cubical container 5cm on a side is 375g, find weight of the liquid in a cubical container 4 cm on a side?

OpenStudy (anonymous):

Assuming the density of the liquid remains constant. let D = density, M = mass and V = volume. \[D = \frac{ M }{ V } =\frac{ 375g }{ 5*5*5 } = \frac{ M_2 }{ 4*4*4 }\]

OpenStudy (anonymous):

cross multiplying will get you the answer

OpenStudy (anonymous):

I came up with 192 g. is that right? im not sure

OpenStudy (anonymous):

it is right

OpenStudy (anonymous):

thank you....I was having difficulties solving this one

OpenStudy (anonymous):

glad i could help :)